OpenStudy (anonymous):

Estimate the quantity in terms of powers of ten 4^50

OpenStudy (llort):

1.2676*10^30

OpenStudy (anonymous):

one way to estimate is to use the fact that 2^10 ~ 1000. So 4^50 = 2^100 = 1000^10, etc.

myininaya (myininaya):

ok 4^1=4 4^2=16 so we need to choose number btw 1 and 2 and itshould be closer to 2 but not too close so we can say 4^1.65 is approximately 10 4^50=approx. 4^(1.65*30)=(4^1.65)^30=10^30